• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name vasulk

    Log entry time 00:11:05 on June15,2002

    Entry number 86035

    keyword=Swing shift summary

    16:00 Begin shift at Kinematic Q2D_c20
    16:10 MCC takes beam to attempt to give us 100 uA
    17:30 Update from MCC, they may get us beam in 30 minutes
    17:40 Moved target to LD2
    17:50 Security Guard came by to alert to a fire alarm that has gone off in Building #99. Someone is coming to check things out. Problem is the heat detector in the beam switch yard.

    18:20 MCC is sending us pulsed beam
    18:30 MCC is ready to send CW beam. Ask for raster on.
    18:50 Thunderstorm. Beam gone.
    18:55 MCC calls. They are having RF problems due to the storm. Time estimate unknown
    19:50 Update from MCC. They can't clear the FSD and are having problems locating the problem.
    21:25 Noticed that Left Q1 tripped. The Left Q1 GUI had two faults: He Lv1 Fault#1 and Fault#2. We left the magnet ramp down to zero. Reset the magnet, and it is ramping back up to the set momentum.
    21:55 Update from MCC. They found that a FSD card needed to be replaced, after rebooting an IOC. They estimate ~ 20 minutes.
    22:35 MCC called they are ready to send us tuned beam.
    23:00 MCC has a minor magnet problem, which they are fixing
    23:15 MCC will send us CW
    23:20 102 uA into Hall A, LD2 target with raster on
    23:25 MCC begins set up for BCM calibration
    23:45 MCC cannot deliver high current beam due to trips. Decide to do Harp scan and start production running. Next shift will address the issue of high current



    Run Summary:

    Run # 1489, pointing run
    Run # 1490, junk need to change prescale factors
    Run # 1491, BCM calibration, beam unstable junk
    Run # 1492, Harp Scan (Run # 1786)